Q5) Use shell method find the volume of the solid generated by revolving the curves y = x2, y = 2 - x2, about y axis
Answer #1

Sola Volume of shell v=2T 8x (2-2²-2²) dx intersection between x =2-x² = x=11. on one side a o to 1 ata sen face of elementar

!! PLEASE VOTE THUMBS UP !!